OpenStudy (anonymous):

The force F of the wind on a flat surface positioned at a right angle to the direction of the wind varies jointly with both the area A of the surface and the square of the speed v of the wind. A wind of 30 miles per hour blowing on a window that that has a surface area of 20 square feet has a force of 150 pounds. What is the force on a window that has a surface area of 12 square feet caused by a wind of 50 miles per hour?

OpenStudy (anonymous):

isnt the formula y=kxz??

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

but isnt it jointly

jimthompson5910 (jim_thompson5910):

"The force F of the wind on a flat surface positioned at a right angle to the direction of the wind varies jointly with both the area A of the surface and the square of the speed v of the wind. " all translates to F = kAv^2

jimthompson5910 (jim_thompson5910):

the formula you wrote is very similar, it's just using different variables

OpenStudy (anonymous):

oh ok go on

jimthompson5910 (jim_thompson5910):

"A wind of 30 miles per hour blowing on a window that that has a surface area of 20 square feet has a force of 150 pounds" means what?

jimthompson5910 (jim_thompson5910):

in terms of the variables given

OpenStudy (anonymous):

30=v,20=a

jimthompson5910 (jim_thompson5910):

Good, v = 30 and A =20, what is F in that sentence?

OpenStudy (anonymous):

150

jimthompson5910 (jim_thompson5910):

so we know that when A = 20 and v = 30, then F = 150 So F = kAv^2 becomes 150 = k*20*(30)^2

jimthompson5910 (jim_thompson5910):

Solve for k

OpenStudy (anonymous):

i dont get a logical answer

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

k=1/120

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so F = kAv^2 becomes F = (1/120)Av^2 which is the same as F = (Av^2)/(120)

jimthompson5910 (jim_thompson5910):

Last part: What is the force on a window that has a surface area of 12 square feet caused by a wind of 50 miles per hour? So you just plug A = 12 and v = 50 into that last equation to find F

OpenStudy (anonymous):

250lbs

OpenStudy (anonymous):

could you help me with the other ones i sent you??

jimthompson5910 (jim_thompson5910):

good, I'm getting that as well

jimthompson5910 (jim_thompson5910):

post them as questions please

OpenStudy (anonymous):

ill post them

jimthompson5910 (jim_thompson5910):

ok